# Nyström Kernel Stein Discrepancy

### v1

@@ 1 Introduction
Estimators for Langevin Stein operator-based KSD exist. But, the classical U-statistic- [Liu et al., 2016] and V-statistic-based [Chwialkowski et al., 2016] estimators have a runtime complexity that scales quadratically with the number of samples of the sampling distribution, which limits their deployment to large-scale settings. To address this bottleneck, Chwialkowski et al. [2016] introduced a linear-time statistic that suffers from low statistical power compared to its quadratic-time counterpart. Jitkrittum et al. [2017] proposed the finite set Stein discrepancy (FSSD), a linear-time approach that replaces the RKHS-norm by the $L_{2}$-norm approximated by sampling; the sampling can either be random (FSSD-rand) or optimized w.r.t. a power proxy (FSSD-opt). Another approach [Huggins and Mackey, 2018] is employing the random Fourier feature (RFF; Rahimi and Recht 2007) method to accelerate the KSD estimation. However, it is known [Chwialkowski et al., 2015, Proposition 1] that the resulting statistic fails to distinguish a large class of measures. Huggins and Mackey [2018] generalize the idea of replacing the RKHS-norm by going from $L_{2}$-norms to $L_{p}$ ones, to obtain feature Stein discrepancies. They present an efficient approximation, random feature Stein discrepancies (RFSD), which is a near-linear time estimator. However, successful deployment of the method depends on a good choice of parameters, which, while the authors provide guidelines, can be challenging to select and tune in practice.
@@ 2 Notations
Let $\left(\mathcal{X},\tau_{\mathcal{X}}\right)$ be a topological space and $\mathcal{B}\left(\tau_{\mathcal{X}}\right)$ the corresponding Borel $\sigma$-algebra. Probability measures in this article are meant w.r.t. the measurable space $\left(\mathcal{X},\mathcal{B}\left(\tau_{\mathcal{X}}\right)\right)$ and are written as $\mathcal{M}_{1}^{+}\left(\mathcal{X}\right)$; for instance, the set of Borel probability measures on $\mathbb{R}^{d}$ is $\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$. The $n$-fold product measure of $\mathbb{P}\in\mathcal{M}_{1}^{+}\left(\mathcal{X}\right)$ is denoted by $\mathbb{P}^{n}\in\mathcal{M}_{1}^{+}\left(\mathcal{X}^{n}\right)$. For a sequence of real-valued random variables $X_{n}$ and a sequence of positive $r_{n}$-s, $X_{n}=\mathcal{O}_{\mathbb{P}}(r_{n})$ means that $\frac{X_{n}}{r_{n}}$ is bounded in probability. The unit ball in a Hilbert space $\mathcal{H}$ is denoted by $B(\mathcal{H})=\left\{f\in\mathcal{H}\mid\left\|f\right\|_{\mathcal{H}}\leq 1\right\}$. The reproducing kernel Hilbert space with $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ as the reproducing kernel is denoted by $\mathcal{H}_{k}$. Throughout the paper, $k$ is assumed to be measurable and $\mathcal{H}_{k}$ to be separable; for instance, a continuous kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ implies the latter property [Steinwart and Christmann, 2008, Lemma 4.33]. The canonical feature map $\phi_{k}:\mathbb{R}^{d}\to\mathcal{H}_{k}$ is defined as $\mathbf{x}\mapsto k(\cdot,\mathbf{x})$; with this feature map $k(\mathbf{x},\mathbf{y})=\left\langle{k(\cdot,\mathbf{x}),k(\cdot,\mathbf{y})}\right\rangle_{\mathcal{H}_{k}}=\left\langle{\phi_{k}(\mathbf{x}),\phi_{k}(\mathbf{y})}\right\rangle_{\mathcal{H}_{k}}$ for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$. The Gram matrix associated with the observations $\left(\mathbf{x}_{i}\right)_{i=1}^{n}\in\left(\mathbb{R}^{d}\right)^{n}$ and kernel $k$ is $\mathbf{K}_{k,n,n}=\left[k(\mathbf{x}_{i},\mathbf{x}_{j})\right]_{i,j=1}^{n}\in\mathbb{R}^{n\times n}$. Given a closed linear subspace $U\subseteq\mathcal{H}_{k}$, the (orthogonal) projection of $h\in\mathcal{H}_{k}$ on $U$ is denoted by $P_{U}h\in U$; $u=P_{U}h$ is the unique vector such that $h-u\perp U$. For any $u\in U$, $\left\|{h-P_{U}h}\right\|_{\mathcal{H}_{k}}\leq\left\|{h-u}\right\|_{\mathcal{H}_{k}}$, that is, $P_{U}h$ is the closest element in $U$ to $h$. A linear operator $A:\mathcal{H}_{k}\to\mathcal{H}_{k}$ is called bounded if $\left\|{A}\right\|_{\mathrm{op}}:=\sup_{\left\|{h}\right\|_{\mathcal{H}_{k}}=1}\left\|{Ah}\right\|_{\mathcal{H}_{k}}<\infty$; the set of $\mathcal{H}_{k}\rightarrow\mathcal{H}_{k}$ bounded linear operators is denoted by $\mathcal{L}(\mathcal{H}_{k})$. An $A\in\mathcal{L}(\mathcal{H}_{k})$ is called positive (shortly $A\geq 0$) if it is self-adjoint ($A^{*}=A$, where $A^{*}\in\mathcal{L}(\mathcal{H}_{k})$ is defined by $\langle Af,g\rangle_{\mathcal{H}_{k}}=\langle f,A^{*}g\rangle_{\mathcal{H}_{k}}$ for all $f,g\in\mathcal{H}_{k}$), and $\langle Ah,h\rangle_{\mathcal{H}_{k}}\geq 0$ for all $h\in\mathcal{H}_{k}$. If $A\geq 0$, then there exists a unique $B\geq 0$ such that $B^{2}=A$; we write $B=A^{\frac{1}{2}}$ and call $B$ the square root of $A$. An $A\in\mathcal{L}(\mathcal{H}_{k})$ is called trace-class if $\sum_{i\in I}\langle(A^{*}A)^{\frac{1}{2}}e_{i},e_{i}\rangle_{\mathcal{H}_{k}}<\infty$ for some countable orthonormal basis (ONB) $(e_{i})_{i\in I}$ of $\mathcal{H}_{k}$, and in this case $\operatorname{tr}(A):=\sum_{i\in I}\langle Ae_{i},e_{i}\rangle_{\mathcal{H}_{k}}<\infty$. [^1] The trace-class property and the value of $\operatorname{tr}(A)$ is independent of the specific ONB chosen. The separability of $\mathcal{H}_{k}$ implies the existence of a countable ONB in it. For a self-adjoint trace-class operator $A$ with eigenvalues $(\lambda_{i})_{i\in I}$, $\operatorname{tr}(A)=\sum_{i\in I}\lambda_{i}$. An operator $A\in\mathcal{L}(\mathcal{H}_{k})$ is called compact if $\overline{\{Ah\,|\,h\in B(\mathcal{H}_{k})\}}$ is compact, where $\overline{\cdot}$ denotes the closure. A trace class operator is compact, and a compact positive operator $A$ has largest eigenvalue $\left\|{A}\right\|_{\mathrm{op}}$. For any $A\in\mathcal{L}(\mathcal{H}_{k})$, it holds that $\left\|{A^{*}A}\right\|_{\mathrm{op}}=\left\|{A}\right\|_{\mathrm{op}}^{2}$ (which is called the $C^{*}$ property).
@@
exists if $\int_{\mathcal{X}}\left\|{\phi_{k}(\mathbf{x})}\right\|_{\mathcal{H}_{k}}^{2}\mathrm{d}\mathbb{P}(\mathbf{x})<\infty$; $C_{\mathbb{P},k}$ is a positive trace-class operator. We define $C_{\mathbb{P},k,\lambda}=C_{\mathbb{P},k}+\lambda I$, where $I$ denotes the identity operator and $\lambda>0$. The effective dimension of $\mathbb{P}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ is defined as $\mathcal{N}_{\mathbb{P},k}(\lambda):=\operatorname{tr}\left(C_{\mathbb{P},k}C_{\mathbb{P},k,\lambda}^{-1}\right)\leq\frac{\operatorname{tr}\left(C_{\mathbb{P},k}\right)}{\lambda}$. [^2] This inequality is implied by $\operatorname{tr}\left(C_{\mathbb{P},k}C_{\mathbb{P},k,\lambda}^{-1}\right)=\sum_{i\in I}\frac{\lambda_{i}}{\lambda_{i}+\lambda}\leq\frac{1}{\lambda}\sum_{i\in I}\lambda_{i}=\frac{\operatorname{tr}\left(C_{\mathbb{P},k}\right)}{\lambda}$, where $(\lambda_{i})_{i\in I}$ denote the eigenvalues of $C_{\mathbb{P},k}$. With $p\geq 1$ and a real-valued random variable $X:\left(\Omega,\mathcal{A},\mathbb{P}\right)\rightarrow\left(\mathbb{R},\mathcal{B}(\tau_{\mathbb{R}})\right)$, where $\mathcal{B}(\tau_{\mathbb{R}})$ denotes the Borel $\sigma$-field on $\mathbb{R}$, let $\left\|X\right\|_{L_{p}(\mathbb{P})}=\left[\int_{\Omega}|X(\omega)|^{p}\mathrm{d}\mathbb{P}(\omega)\right]^{\frac{1}{p}}$. For $p\in\{1,2\}$, let $\psi_{p}(u)=e^{u^{p}}-1$ and $\left\|{X}\right\|_{\psi_{p}}:=\inf\left\{C>0\mid\mathbb{E}_{X\sim\mathbb{P}}\psi_{p}\left(\frac{|X|}{C}\right)\leq 1\right\}$. A real-valued random variable $X\sim\mathbb{P}\in\mathcal{M}_{1}^{+}(\mathbb{R})$ is called sub-exponential if $\left\|{X}\right\|_{\psi_{1}}<\infty$ and sub-Gaussian if $\left\|{X}\right\|_{\psi_{2}}<\infty$. In the following, we specialize Definition 2 by Koltchinskii and Lounici [2017] stated for Banach spaces to (reproducing kernel) Hilbert spaces by using the Riesz representation theorem. A centered random variable $X\sim\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathcal{H}_{k}\right)$ taking values in an RKHS $\mathcal{H}_{k}$ is called sub-Gaussian iff. there exists a universal constant $C>0$ such that
@@ 3 Problem formulation
notice that $\bm{\xi}_{p}(\mathbf{x})$ and $h_{p}(\cdot,\mathbf{x})$ map to different feature spaces ($\mathcal{H}_{k}^{d}$ and $\mathcal{H}_{h_{p}}$, respectively) but yield the same kernel $h_{p}$. The kernel Stein discrepancy (KSD) [Chwialkowski et al., 2016, Liu et al., 2016] then is defined as
@@ 4.1 The Nyström-KSD estimator
    Runtime complexity. The computation of (12) consists of calculating $\bm{\beta}_{p}$, pseudo-inverting $\bm{K}_{h_{p},m,m}$, and obtaining the quadratic form $\bm{\beta}_{p}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m,m}^{-}\bm{\beta}_{p}$. The calculation of $\bm{\beta}_{p}$ requires $\mathcal{O}(mn)$ operations, due to the multiplication of an $m\times n$ matrix with a vector of length $n$. Inverting the $m\times m$ matrix $\bm{K}_{h_{p},m,m}$ costs $\mathcal{O}(m^{3})$, [^4] Although faster algorithms for matrix inversion exist, e.g., Strassen’s algorithm, we consider the runtime that one typically encounters in practice. dominating the cost of $\mathcal{O}(m^{2})$ needed for the computation of $\bm{K}_{h_{p},m,m}$. The quadratic form $\bm{\beta}_{p}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m,m}^{-}\bm{\beta}_{p}$ has a computational cost of $\mathcal{O}\left(m^{2}\right)$. Hence, (12) can be computed in $\mathcal{O}\left(mn+m^{3}\right)$, which means that for $m=o\left(n^{2/3}\right)$, our proposed Nyström-KSD estimator guarantees an asymptotic speedup.

### v2

@@ 1 Introduction
Estimators for Langevin Stein operator-based KSD exist. But, the classical U-statistic- [Liu et al., 2016] and V-statistic-based [Chwialkowski et al., 2016] estimators have a runtime complexity that scales quadratically with the number of samples of the sampling distribution, which limits their deployment to large-scale settings. To address this bottleneck, Chwialkowski et al. [2016] introduced a linear-time statistic that suffers from low statistical power compared to its quadratic-time counterpart. Jitkrittum et al. [2017] proposed the finite set Stein discrepancy (FSSD), a linear-time approach that replaces the RKHS-norm by the $L_{2}$-norm approximated by sampling; the sampling can either be random (FSSD-rand) or optimized w.r.t. a power proxy (FSSD-opt). Another approach [Huggins and Mackey, 2018] is employing the random Fourier feature (RFF; Rahimi and Recht 2007) method to accelerate the KSD estimation. However, it is known [Chwialkowski et al., 2015, Proposition 1] that the resulting statistic fails to distinguish a large class of measures. Huggins and Mackey [2018] generalize the idea of replacing the RKHS-norm by going from $L_{2}$-norms to $L_{p}$ ones, to obtain feature Stein discrepancies. They present an efficient approximation, random feature Stein discrepancies (RFSD), which is a (near-)linear time estimator. However, successful deployment of the method depends on a good choice of parameters, which, while the authors provide guidelines, can be challenging to select and tune in practice.
@@ 2 Notations
Let $\left(\mathcal{X},\tau_{\mathcal{X}}\right)$ be a topological space and $\mathcal{B}\left(\tau_{\mathcal{X}}\right)$ the corresponding Borel $\sigma$-algebra. Probability measures in this article are meant w.r.t. the measurable space $\left(\mathcal{X},\mathcal{B}\left(\tau_{\mathcal{X}}\right)\right)$ and are written as $\mathcal{M}_{1}^{+}\left(\mathcal{X}\right)$; for instance, the set of Borel probability measures on $\mathbb{R}^{d}$ is $\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$. The $n$-fold product measure of $\mathbb{P}\in\mathcal{M}_{1}^{+}\left(\mathcal{X}\right)$ is denoted by $\mathbb{P}^{n}\in\mathcal{M}_{1}^{+}\left(\mathcal{X}^{n}\right)$. The product of $\mathbb{P}_{1}\in\mathcal{M}_{1}^{+}(\mathcal{X}_{1})$ and $\mathbb{P}_{2}\in\mathcal{M}_{1}^{+}(\mathcal{X}_{2})$ is written as $\mathbb{P}_{1}\otimes\mathbb{P}_{2}$ ($\in\mathcal{M}_{1}^{+}(\mathcal{X}_{1}\times\mathcal{X}_{2})$), where $\left(\mathcal{X}_{1},\tau_{\mathcal{X}_{1}}\right)$ and $\left(\mathcal{X}_{2},\tau_{\mathcal{X}_{2}}\right)$ are topological spaces. For a sequence of i.i.d. real-valued random variables $X_{n}\sim\mathbb{P}\in\mathcal{M}_{1}^{+}(\mathbb{R})$ and a sequence of positive $r_{n}$-s, $X_{n}=\mathcal{O}_{\mathbb{P}}(r_{n})$ means that $\frac{X_{n}}{r_{n}}$ is bounded in probability. The unit ball in a Hilbert space $\mathcal{H}$ is denoted by $B(\mathcal{H})=\left\{f\in\mathcal{H}\mid\left\|f\right\|_{\mathcal{H}}\leq 1\right\}$. The reproducing kernel Hilbert space with $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ as the reproducing kernel is denoted by $\mathcal{H}_{k}$. Throughout the paper, $k$ is assumed to be measurable and $\mathcal{H}_{k}$ to be separable; for instance, a continuous kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ implies the latter property [Steinwart and Christmann, 2008, Lemma 4.33]. The canonical feature map $\phi_{k}:\mathbb{R}^{d}\to\mathcal{H}_{k}$ is defined as $\mathbf{x}\mapsto k(\cdot,\mathbf{x})$; with this feature map $k(\mathbf{x},\mathbf{y})=\left\langle{k(\cdot,\mathbf{x}),k(\cdot,\mathbf{y})}\right\rangle_{\mathcal{H}_{k}}=\left\langle{\phi_{k}(\mathbf{x}),\phi_{k}(\mathbf{y})}\right\rangle_{\mathcal{H}_{k}}$ for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$. The Gram matrix associated with the observations $\left(\mathbf{x}_{i}\right)_{i=1}^{n}\in\left(\mathbb{R}^{d}\right)^{n}$ and kernel $k$ is $\mathbf{K}_{k,n,n}=\left[k(\mathbf{x}_{i},\mathbf{x}_{j})\right]_{i,j=1}^{n}\in\mathbb{R}^{n\times n}$. Given a closed linear subspace $U\subseteq\mathcal{H}_{k}$, the (orthogonal) projection of $h\in\mathcal{H}_{k}$ on $U$ is denoted by $P_{U}h\in U$; $u=P_{U}h$ is the unique vector such that $h-u\perp U$. For any $u\in U$, $\left\|{h-P_{U}h}\right\|_{\mathcal{H}_{k}}\leq\left\|{h-u}\right\|_{\mathcal{H}_{k}}$, that is, $P_{U}h$ is the closest element in $U$ to $h$. A linear operator $A:\mathcal{H}_{k}\to\mathcal{H}_{k}$ is called bounded if $\left\|{A}\right\|_{\mathrm{op}}:=\sup_{\left\|{h}\right\|_{\mathcal{H}_{k}}=1}\left\|{Ah}\right\|_{\mathcal{H}_{k}}<\infty$; the set of $\mathcal{H}_{k}\rightarrow\mathcal{H}_{k}$ bounded linear operators is denoted by $\mathcal{L}(\mathcal{H}_{k})$. An $A\in\mathcal{L}(\mathcal{H}_{k})$ is called positive (shortly $A\geq 0$) if it is self-adjoint ($A^{*}=A$, where $A^{*}\in\mathcal{L}(\mathcal{H}_{k})$ is defined by $\langle Af,g\rangle_{\mathcal{H}_{k}}=\langle f,A^{*}g\rangle_{\mathcal{H}_{k}}$ for all $f,g\in\mathcal{H}_{k}$), and $\langle Ah,h\rangle_{\mathcal{H}_{k}}\geq 0$ for all $h\in\mathcal{H}_{k}$. If $A\geq 0$, then there exists a unique $B\geq 0$ such that $B^{2}=A$; we write $B=A^{\frac{1}{2}}$ and call $B$ the square root of $A$. An $A\in\mathcal{L}(\mathcal{H}_{k})$ is called trace-class if $\sum_{i\in I}\langle(A^{*}A)^{\frac{1}{2}}e_{i},e_{i}\rangle_{\mathcal{H}_{k}}<\infty$ for some countable orthonormal basis (ONB) $(e_{i})_{i\in I}$ of $\mathcal{H}_{k}$, and in this case $\operatorname{tr}(A):=\sum_{i\in I}\langle Ae_{i},e_{i}\rangle_{\mathcal{H}_{k}}<\infty$. [^1] The trace-class property and the value of $\operatorname{tr}(A)$ is independent of the specific ONB chosen. The separability of $\mathcal{H}_{k}$ implies the existence of a countable ONB in it. For a self-adjoint trace-class operator $A$ with eigenvalues $(\lambda_{i})_{i\in I}$, $\operatorname{tr}(A)=\sum_{i\in I}\lambda_{i}$. An operator $A\in\mathcal{L}(\mathcal{H}_{k})$ is called compact if $\overline{\{Ah\,|\,h\in B(\mathcal{H}_{k})\}}$ is compact, where $\overline{\cdot}$ denotes the closure. A trace class operator is compact, and a compact positive operator $A$ has largest eigenvalue $\left\|{A}\right\|_{\mathrm{op}}$. For any $A\in\mathcal{L}(\mathcal{H}_{k})$, it holds that $\left\|{A^{*}A}\right\|_{\mathrm{op}}=\left\|{A}\right\|_{\mathrm{op}}^{2}$ (which is called the $C^{*}$ property).
@@
exists if $\int_{\mathcal{X}}\left\|{\phi_{k}(\mathbf{x})}\right\|_{\mathcal{H}_{k}}^{2}\mathrm{d}\mathbb{P}(\mathbf{x})<\infty$; $C_{\mathbb{P},k}$ is a positive trace-class operator. We define $C_{\mathbb{P},k,\lambda}=C_{\mathbb{P},k}+\lambda I$, where $I$ denotes the identity operator and $\lambda>0$. The effective dimension of $\mathbb{P}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ is defined as $\mathcal{N}_{\mathbb{P},k}(\lambda):=\operatorname{tr}\left(C_{\mathbb{P},k}C_{\mathbb{P},k,\lambda}^{-1}\right)\leq\frac{\operatorname{tr}\left(C_{\mathbb{P},k}\right)}{\lambda}$. [^2] This inequality is implied by $\operatorname{tr}\left(C_{\mathbb{P},k}C_{\mathbb{P},k,\lambda}^{-1}\right)=\sum_{i\in I}\frac{\lambda_{i}}{\lambda_{i}+\lambda}\leq\frac{1}{\lambda}\sum_{i\in I}\lambda_{i}=\frac{\operatorname{tr}\left(C_{\mathbb{P},k}\right)}{\lambda}$, where $(\lambda_{i})_{i\in I}$ denote the eigenvalues of $C_{\mathbb{P},k}$. With $r\geq 1$ and a real-valued random variable $X:\left(\Omega,\mathcal{A},\mathbb{P}\right)\rightarrow\left(\mathbb{R},\mathcal{B}(\tau_{\mathbb{R}})\right)$, where $\mathcal{B}(\tau_{\mathbb{R}})$ denotes the Borel $\sigma$-field on $\mathbb{R}$, let $\left\|X\right\|_{L_{r}(\mathbb{P})}=\left[\int_{\Omega}|X(\omega)|^{r}\mathrm{d}\mathbb{P}(\omega)\right]^{\frac{1}{r}}$. For $r\in\{1,2\}$, let $\psi_{r}(u)=e^{u^{r}}-1$ and $\left\|{X}\right\|_{\psi_{r}}:=\inf\left\{C>0\mid\mathbb{E}_{X\sim\mathbb{P}}\psi_{r}\left(\frac{|X|}{C}\right)\leq 1\right\}$. A real-valued random variable $X\sim\mathbb{P}\in\mathcal{M}_{1}^{+}(\mathbb{R})$ is called sub-exponential if $\left\|{X}\right\|_{\psi_{1}}<\infty$ and sub-Gaussian if $\left\|{X}\right\|_{\psi_{2}}<\infty$. In the following, we specialize Definition 2 by Koltchinskii and Lounici [2017] stated for Banach spaces to (reproducing kernel) Hilbert spaces by using the Riesz representation theorem. A centered random variable $X\sim\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathcal{H}_{k}\right)$ taking values in an RKHS $\mathcal{H}_{k}$ is called sub-Gaussian iff. there exists a universal constant $C>0$ such that
@@ 3 Problem formulation
notice that $\bm{\xi}_{p}(\mathbf{x})$ and $h_{p}(\cdot,\mathbf{x})$ map to different feature spaces ($\mathcal{H}_{k}^{d}$ and $\mathcal{H}_{h_{p}}$, respectively) but yield the same kernel $h_{p}$. Throughout this work, we assume the measurability of $h_{p}$ and the separability of $\mathcal{H}_{h_{p}}$; this is satisfied, for example, if $k$ is continuously differentiable.
The kernel Stein discrepancy (KSD) [Chwialkowski et al., 2016, Liu et al., 2016] then is defined as
@@ 4.1 The Nyström-KSD estimator
    Runtime complexity. The computation of (12) consists of calculating $\bm{\beta}_{p}$, pseudo-inverting $\bm{K}_{h_{p},m,m}$, and obtaining the quadratic form $\bm{\beta}_{p}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m,m}^{-}\bm{\beta}_{p}$. The calculation of $\bm{\beta}_{p}$ requires $\mathcal{O}(mn)$ operations, due to the multiplication of an $m\times n$ matrix with a vector of length $n$. Inverting the $m\times m$ matrix $\bm{K}_{h_{p},m,m}$ costs $\mathcal{O}(m^{3})$, [^4] Although faster algorithms for (pseudo) matrix inversion exist, we consider the runtime that one typically encounters in practice. dominating the cost of $\mathcal{O}(m^{2})$ needed for the computation of $\bm{K}_{h_{p},m,m}$. The quadratic form $\bm{\beta}_{p}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m,m}^{-}\bm{\beta}_{p}$ has a computational cost of $\mathcal{O}\left(m^{2}\right)$. Hence, (12) can be computed in $\mathcal{O}\left(mn+m^{3}\right)$, which means that for $m=o\left(n^{2/3}\right)$, our proposed Nyström-KSD estimator guarantees an asymptotic speedup.